# LoXR: Performance Evaluation of Locally Executing LLMs on XR Devices

### v1

@@ Abstract
The deployment of large language models (LLMs) on extended reality (XR) devices has great potential to advance the field of human-AI interaction. In case of direct, on-device model inference, selecting the appropriate model and device for specific tasks remains challenging. In this paper, we deploy 17 LLMs across four XR devices—Magic Leap 2, Meta Quest 3, Vivo X100s Pro, and Apple Vision Pro—and conduct a comprehensive evaluation. We devise an experimental setup and evaluate performance on four key metrics: performance consistency, processing speed, memory usage, and battery consumption. For each of the 68 model-device pairs, we assess performance under varying string lengths, batch sizes, and thread counts, analyzing the tradeoffs for real-time XR applications. We finally propose a unified evaluation method based on the Pareto Optimality theory to select the optimal device-model pairs from the quality and speed objectives. We believe our findings offer valuable insight to guide future optimization efforts for LLM deployment on XR devices. Our evaluation method can be followed as standard groundwork for further research and development in this emerging field. All supplemental materials are available at nanovis.org/Loxr.html.
@@ 1 Introduction
Since the 2022 release of OpenAI’s ChatGPT interface [1], relying on the GPT-3.5 Large Language Model (LLM), we have witnessed the disruptive and transformative effects of LLMs. These models are capable of describing a wide variety of topics, respond at various levels of abstraction, and communicate effectively in multiple languages. They have proven capable of providing users with accurate and contextually appropriate responses. LLMs have quickly found applications in tasks such as spelling and grammar correction [2], generating text on specified topics [3], integration into automated chatbot services, and even generating source code from loosely defined software specifications [4].
Research on language models, and on their multimodal variants integrating language and vision or other technologies has recently experienced rapid growth. For instance, in computer vision, language models are combined with visual signals to achieve tasks such as verbal scene description and even open-world scenegraph generation [5]. These technologies enable detailed interpretation of everyday objects, inference of relationships among them, and estimates of physical properties like size, weight, distance, and speed. In user interaction and visualization research, LLMs serve as verbal interfaces to control software functionality or adjust visualization parameters [6, 7]. Through prompt engineering or fine-tuning, loosely defined text can be translated into specific commands that execute desired actions within a system, supported by language model APIs. The capabilities of language models continue to improve significantly from one version to the next. Yet, this comes at the cost of a significant growth in size of the most advanced models. As an example, models released or operated by leading companies such as OpenAI, Meta, Microsoft, and Google now reach the scale of trillions of trainable parameters [1]. Due to their size and complexity, the training and inference of these models are limited to dedicated data centers equipped with considerable computational and memory resources.
@@
Another disruptive application of LLMs involves their use in wearable technology to assist users in various environments. For instance, Meta AI is being integrated into Meta’s virtual reality hardware, where images captured by the device are streamed over the network to Meta’s servers for model inference, enabling image interpretation or prompt response tasks. [^1] https://www.meta.com/blog/quest/meta-ai-on-meta-quest-3/ The overarching vision is to develop assistive hardware that is as lightweight as sunglasses but capable of aiding users in a wide range of tasks by understanding the scene they are viewing.
While such technological advances have the potential to empower users in unprecedented ways, three main drawbacks must be considered due to the heavy reliance on network connectivity and cloud-based data processing. First, users will need to maintain extremely high network connectivity to ensure that remote services can perform assistive tasks without excessive latency, which would degrade the user experience. Second, data privacy poses a significant challenge. Many environments, such as medical settings and various industries, operate under strict data privacy regulations that prohibit the transmission of sensitive information to remote servers, precluding the use of these services in such contexts. Third, the reliance on subscription models for these assistive services introduces ongoing costs for users, requiring them to periodically purchase licenses, which can accumulate into significant long-term costs.
Therefore, there is dedicated research effort focused on developing autonomous AI assistive technologies that can perform inference directly on a device. To this goal, several smaller LLMs have been released by Meta, Mistral AI, Microsoft, and Google [8]. These models are designed to potentially operate on mobile devices or extended reality hardware and are typically around billions of trainable parameters in size. Some LLMs with fewer than a billion trainable parameters have shown very good results [8].
@@
Local execution of LLMs on XR devices is expected to become increasingly necessary for a wide range of applications, particularly those that require real-time processing and device autonomy. However, selecting the optimal device and model for a specific application is a complex decision: device specifications and model documentation alone are not sufficient for making it, as the performance of most LLMs running on various hardware components remains largely undocumented. Furthermore, the hardware configurations of XR devices (such as CPU, memory capacity, and thermal management) and the architectures of the LLMs can significantly influence the results. Therefore, defining standardized evaluation criteria that address these complexities and ensures fairness across tests is also a research challenge. To the best of our knowledge, no comprehensive study currently provides clear guidelines for such evaluations.
@@
To address these challenges, we present LoXR, a comprehensive study deploying LLMs on XR, focusing on performance evaluation across various metrics. Considering the multiple challenges and factors affecting the evaluation, we first identify key evaluation metrics, and then design an experimental setup to minimize potential bias. Specifically, we deploy 17 LLMs across four XR devices—Magic Leap 2, Meta Quest 3, Vivo X100s Pro [^2] Vivo X100s Pro is strictly speaking not an XR device. We include it in our comparison because smartphones are commonly used as XR devices, and this device features competitive hardware specifications in the market. , and Apple Vision Pro—and conduct a thorough evaluation in terms of performance consistency i.e., stability over time, processing speed, memory usage, and battery consumption. This results in 68 model-device pairs. A key challenge is the effect of varying string lengths, batch sizes, thread counts, and background processes. To account for string length variation, we conducted experiments using different lengths. To mitigate hidden factors like background processes, each experiment was repeated five times, and average values were calculated. To provide task-specific analysis, experiments were conducted across all LLMs tasks: Prompt Processing (PP) for evaluating performance with varying prompt lengths, Token Generation (TG) for varying generated token set sizes, plus Batch Test (BT) and Thread Test (TT) for assessing parallelism and concurrency efficiency. We believe our experimental protocol and results will serve as a foundation for further research and development in this emerging field.
In summary, this paper contributes:
@@
  • An evaluation approach and an experimental setup for assessing the performance of LLMs on XR devices, that can be used as standard guidelines for future research.
@@
  • The deployment of 17 LLMs on four XR devices, and a comprehensive evaluation, resulting in 68 model-device pairs, across five critical metrics: consistency, processing speed, concurrency and parallelism, memory usage, and battery consumption.
@@ 2 Related Work
We review related work with respect to two major research directions. The first research direction is exploring the use of LLMs on XR devices, and the second research direction explores how to make LLMs directly deployable on low-resource hardware.
@@ 2.1 LLM-powered XR Applications
Owing to the powerful semantic understanding and extensive general knowledge of LLMs, numerous studies have explored their application in assisting various tasks within XR scenarios.
Torre et al. [9] introduced LLMR, a framework that leverages LLMs for the real-time creation and modification of interactive mixed reality experiences, enabling tasks such as generating new content or editing existing works on VR/AR devices. Jia et al. [6] developed the VOICE framework, which employs a two-layer agent system for conversational interaction and explanation in scientific communication, with a prototype deployable on VR devices. Kurai et al. [10] proposed MagicItem, a tool that allows users with limited programming experience to define object behaviors within the metaverse platform. Zhang et al. [11] introduced OdorAgent, which combines an LLM with a text-image model to automate video-odor matching. Yin et al. [12] identified potential limitations in LLM-based automated systems and proposed the systematic framework Text2VRScene to address them. Chen et al. [13] leveraged the extensive capabilities of LLMs in context perception and text prediction to enhance text entry efficiency by reducing manual keystrokes in VR scenarios. Giunchi et al. [14] developed DreamCodeVR, a tool designed to help users, regardless of coding experience, create basic object behavior in VR environments by translating spoken language into code within an active application. Wan et al. [15] presented an LLM-based AI agent for human-agent interaction in VR, involving GPT-4 to simulate realistic NPC behavior, including context-aware responses, facial expressions, and body gestures.
As they stand, these previous contributions still require the deployed devices to connect to a cloud server for LLM inference, raising concerns about user privacy, latency, costs, and internet access requirements. In our work, we evaluate the local inference performance of various LLMs and propose a prototype capable of using local LLM inference.
@@ 3 Research Methods and Materials
The aim of this study is to deploy LLMs on XR devices and evaluate their performance in different aspects. This section outlines our research methodology, including the key challenges, the LLM deployment approach, and the evaluation metrics and models used in this study.
Problem Statement: In this study, we deploy 17 LLMs, denoted as $\{m_{1},m_{2},\dots,m_{17}\}$ (see Table I), across four XR and mobile devices: $d_{1},d_{2},d_{3},$ and $d_{4}$, representing Magic Leap 2 (ML2), [^3] https://www.magicleap.com/magicleap-2 Meta Quest 3 (MQ3), [^4] https://www.meta.com/quest Vivo X100 Pro (Vivo), [^5] https://www.vivo.com/en and Apple Vision Pro (AVP), [^6] https://www.apple.com/vision-pro respectively. Our goal is to perform a comprehensive evaluation of these models on the four devices to determine how well they handle large-scale language processing tasks in resource-constrained XR environments.
We define our evaluation using six key performance metrics: $p_{0},p_{1},p_{2},p_{3},p_{4},p_{5}$, which represent Model Quality, Performance Consistency, Performance Speed (with respect to string length and prompt length), Parallelism (with respect to thread count and varying batch size), Memory Usage, and Battery Usage, respectively.
Each LLM $m_{i}\in\{m_{1},m_{2},\dots,m_{17}\}$ is tested on each XR device $d_{j}\in\{d_{1},d_{2},d_{3},d_{4}\}$ for evaluation across each performance metric $p_{k}\in\{p_{0},p_{1},p_{2},p_{3},p_{4},p_{5}\}.$ The evaluation can be summarized as $X_{mdk}$, as given in Eq. 1.
$X_{mdk}=\{x_{ijk}\mid i\in\{1,\dots,17\},j\in\{1,\dots,4\},k\in\{0,\dots,5\}\}$ (1)
Here, $x_{ijk}$ represents the performance of model $m_{i}$ on device $d_{j}$ for metric $p_{k}$, forming the basis for cross-device and cross-model analysis.
Fig. 1 presents an overview of our research pipeline, where each rectangular box represents a specific procedure, and the connected ovals indicate their respective outcomes. These are discussed in detail below.
Figure 1: Overview of our research process. The rectangular boxes represent the procedures, while the ovals indicate the outcomes. We deployed 17 LLMs on four XR devices. First, we identified the key factors affecting performance. Based on these factors, we defined our evaluation criteria and conducted the experiments. Finally, we applied 3D Pareto analysis, which provided the Pareto front to highlight the best-performing model-device pairs.
@@ 3.3 Evaluation Approach
Considering the key challenges and our initial investigation (see Section 3.2), we formulate a standardized evaluation approach for our assessments. First, we designed an experimental setup to minimize potential biases and ensure the validity of the results (see Section 4). This setup establishes identical parameters across all devices and defines the test types to be executed.
We developed several guidelines to ensure a comprehensive and standardized evaluation. To account for potential performance variations among LLMs, we selected 17 models from five different series (see Table I) and evaluated them across four devices. Additionally, we conducted Model Quality analysis to provide technical insights into the selected models. These tests are device-independent, and were conducted to understand the details and properties of each model. Next, each model-device pair was subjected to various tests (see Section 3.4).
@@ 3.4 Models and Evaluation Metrics
We use 17 LLMs from five different series (see Table I) in our evaluations. The models were obtained from Hugging Face’s model repository, a widely used platform for accessing pre-trained LLMs [23]. For consistency in naming conventions, we assigned unique IDs ($m_{1}$ to $m_{17}$) to these models.
@@ 4 Experimental Setup
Processing Speed: In this study, we measure the processing speed for each model-device pair $(m_{i},d_{j})$ by recording the time $t_{ijk}$ taken by model $m_{i}$ running on device $d_{j}$ to process a string of length $L$. The instantaneous processing speed, $S_{inst.}$, expressed in tokens per second, is calculated as follows:
$S_{inst.}=\frac{L}{t_{ij}}.$ (2)
However, instead of relying on a single run, we perform 5 runs for each model-device pair to ensure accuracy and measure stability. If the coefficient of variance (CV) across the 5 runs is less than $40\%$, we calculate the 5-Run-Mean Speed (referred to as consistent speed, $S_{c}$) as:
@@
where $S_{\text{inst.}}(k)$ represents the instantaneous speed during the $k^{\text{th}}$ run of each experiment, calculated via Eq. 2. This approach minimizes the impact of factors such as initialization delays, hardware fluctuations, or other environmental inconsistencies, providing a more reliable measure of processing speed.
@@
We conducted the following four types of experiments.

### v2

# AIvaluateXR: An Evaluation Framework for on-Device AI in XR with Benchmarking Results

@@ Abstract
The deployment of large language models (LLMs) on extended reality (XR) devices has great potential to advance the field of human-AI interaction. In case of direct, on-device model inference, selecting the appropriate model and device for specific tasks remains challenging. In this paper, we present AIvaluateXR, a comprehensive evaluation framework for benchmarking LLMs running on XR devices. To demonstrate the framework, we deploy 17 selected LLMs across four XR platforms—Magic Leap 2, Meta Quest 3, Vivo X100s Pro, and Apple Vision Pro—and conduct an extensive evaluation. Our experimental setup measures four key metrics: performance consistency, processing speed, memory usage, and battery consumption. For each of the 68 model-device pairs, we assess performance under varying string lengths, batch sizes, and thread counts, analyzing the tradeoffs for real-time XR applications. We finally propose a unified evaluation method based on the 3D Pareto Optimality theory to select the optimal device-model pairs from the quality and speed objectives. Additionally, we compare the efficiency of on-device LLMs with client-server and cloud-based setups, and evaluate their accuracy on two interactive tasks. We believe our findings offer valuable insight to guide future optimization efforts for LLM deployment on XR devices. Our evaluation method can be followed as standard groundwork for further research and development in this emerging field. The source code and supplementary materials are available at nanovis.org/AIvaluateXR.html.
@@ 1 Introduction
Extended Reality (XR) and Artificial Intelligence (AI) have become increasingly prominent, and their intersection is even more compelling for researchers and real-world applications [1, 2, 3, 4, 5, 6]. Since the 2022 release of OpenAI’s ChatGPT interface [7], relying on the GPT-3.5 Large Language Model (LLM), we have witnessed the disruptive and transformative effects of LLMs. These models are capable of describing a wide variety of topics, respond at various levels of abstraction, and communicate effectively in multiple languages. They have proven capable of providing users with accurate and contextually appropriate responses. In addition to general applications such as grammar and text processing [8, 9], automated chatbot services, and source code generation [10], LLMs are increasingly being utilized in computer vision, human-computer interaction, and visual understanding.
@@
In computer vision, language models are combined with visual signals to perform tasks such as verbal scene description and open-world scene graph generation [11]. In user interaction and visualization research, LLMs serve as verbal interfaces for controlling software functionality or adjusting visualization parameters [12, 13]. Another disruptive application of LLMs involves their use in wearable technology [14] to assist users in various environments. For instance, Meta AI is being integrated into Meta’s virtual reality hardware, where images captured by the device are streamed over the network to Meta’s servers for model inference, enabling image interpretation or prompt response tasks [^1] https://www.meta.com/blog/quest/meta-ai-on-meta-quest-3/. The overarching vision is to develop assistive hardware that is as lightweight as sunglasses but capable of aiding users in a wide range of tasks by understanding the scene they are viewing.
While such technological advances have the potential to empower users in unprecedented ways, the need for on-device AI in XR arises in several scenarios: (i) regions with limited or no internet connectivity (e.g., field training or remote maintenance); (ii) strict security and privacy requirements (e.g., healthcare, defense); (iii) reducing network latency; (iv) enabling autonomous operation as both LLMs and XR hardware continue to advance rapidly; (v) avoiding subscription costs; and (vi) providing the ability to fine-tune models for specific tasks and domains. In this regard, several smaller LLMs have been released by Meta, Mistral AI, Microsoft, and Google [15]. These models are designed to operate efficiently on mobile devices and extended reality hardware, typically consisting of billions of trainable parameters, while some LLMs with fewer than a billion parameters have already demonstrated promising performance [15].
@@
Local execution of LLMs on XR devices is expected to become increasingly necessary for a wide range of applications. However, selecting the optimal device and model for a specific application is a complex decision. Device specifications and model documentation alone are not sufficient for making this decision. Furthermore, the hardware configurations of XR devices (such as CPU, memory capacity, and thermal management) and the architectures of the LLMs can significantly influence the results. Therefore, defining standardized evaluation criteria that address these complexities and ensures fairness across tests is also a research challenge. To the best of our knowledge, no comprehensive study currently provides clear guidelines for such evaluations.
To address these challenges, we present AIvaluateXR, a comprehensive framework for deploying LLMs on XR platforms and systematically evaluating their performance across multiple metrics. AIvaluateXR works in two steps: first, it deploys LLMs on XR devices and evaluates their performance across a range of factors; then, in the second step, it applies Pareto analysis to the results to identify the best-performing model-device pairs. We applied AIvaluateXR to deploy 17 LLMs across four XR devices—Magic Leap 2, Meta Quest 3, Vivo X100s Pro [^2] Vivo X100s Pro is, strictly speaking, not an XR device. We include it in our comparison because smartphones are commonly used as XR platforms, and this device features competitive hardware specifications., and Apple Vision Pro—and conducted a thorough evaluation in terms of performance consistency (i.e., stability over time), processing speed, memory usage, and battery consumption. Finally, we used Pareto analysis to identify the best-performing model-device pairs among the 68 combinations evaluated.
@@
Considering the multiple challenges and factors affecting the evaluation, we first identify key evaluation metrics, and then design an experimental setup to minimize potential bias. A key challenge is the effect of varying string lengths, batch sizes, thread counts, and background processes. To account for string length variation, we conducted experiments using different lengths. To mitigate hidden factors like background processes, each experiment was repeated five times, and average values were calculated. To provide task-specific analysis, experiments were conducted across all LLMs tasks: Prompt Processing (PP) for evaluating performance with varying prompt lengths, Token Generation (TG) for varying generated token set sizes, plus Batch Test (BT) and Thread Test (TT) for assessing parallelism and concurrency efficiency. We believe our experimental protocol and results will serve as a foundation for further research and development in this emerging field. In summary, this paper contributes:
@@
  • An evaluation framework and an experimental setup for assessing the performance of LLMs on XR devices, which can be used as standard guidelines for future research.
@@
  • A comprehensive evaluation of 17 LLMs on four XR devices, resulting in 68 model-device pairs, across four critical metrics: processing speed, memory usage, stability over time, and battery consumption.
  • The deployment of various LLMs on XR devices, along with an analysis of their accuracy for interactive applications using standard query datasets, and a comparison with cloud-based and client-server-based LLMs in terms of efficiency and accuracy.
@@ 2 Related Work
We review related work along two major research directions: LLM-powered XR and on-device LLMs. For further details on AI in XR, see the recent review articles [1, 2, 4, 3].
@@ 2.1 LLM-powered XR Applications
Owing to the powerful semantic understanding and extensive general knowledge of LLMs, numerous studies have explored their application in assisting various tasks within XR scenarios. Recent studies have demonstrated the use of LLMs for natural interaction and conversational interfaces in VR and AR scenarios [16, 17, 18]. LLMs have also been applied for content generation, chat-based assistance, VR education[19], social VR [20], and as intelligent companions in XR applications [21, 22]. In addition to LLMs, vision-language models (VLMs) have been integrated into AR systems to enhance visual and multimodal understanding [14], scene-driven AR storytelling [23, 24], contextual guidance, dynamic content creation, and much more [25].
XaiR [14] integrates Multimodal Large Language Models (MLLMs) with XR devices using a client-server architecture, where computationally intensive MLLM tasks are offloaded to a server while spatial context is handled locally on the XR headset. The system supports real-time multimodal input; however, latency—comprising both network delays and LLM inference time—remains a challenge, often requiring users to wait 4–6 seconds for a response. Torre et al.[26] introduced a similar framework, LLMR, which leverages LLMs for real-time creation and modification of interactive mixed reality experiences, enabling tasks such as generating new content or editing existing elements on VR/AR devices.
Jia et al. [12] developed the VOICE framework, which employs a two-layer agent system for conversational interaction and explanation in scientific communication, with a prototype deployable on VR devices. Kurai et al. [27] proposed MagicItem, a tool that allows users with limited programming experience to define object behaviors within the metaverse platform. Zhang et al. [28] introduced OdorAgent, which combines an LLM with a text-image model to automate video-odor matching. Yin et al. [29] identified potential limitations in LLM-based automated systems and proposed the systematic framework Text2VRScene to address them. Chen et al. [30] leveraged the extensive capabilities of LLMs in context perception and text prediction to enhance text entry efficiency by reducing manual keystrokes in VR scenarios. Giunchi et al. [31] developed DreamCodeVR, a tool designed to help users, regardless of coding experience, create basic object behavior in VR environments by translating spoken language into code within an active application. Wan et al. [32] presented an LLM-based AI agent for human-agent interaction in VR, involving GPT-4 to simulate realistic NPC behavior, including context-aware responses, facial expressions, and body gestures.
XR applications often require 3D content that can be generated using AI. For example, Kawka et al. [33] proposed a method to create 3D objects from textual and image data, along with techniques for surface remeshing and simplification. Other approaches include character animation [34], text-to-3D [35], speech-to-3D generations [36], and image-to-3D reconstruction [37]. Recent work has also explored modeling human avatars and motion [38], as well as spatial understanding and object recommendation [39]. Although these systems typically rely on cloud-based models or client-server architectures rather than performing inference directly on the device, they demonstrate a clear need for AI in XR. This dependence raises concerns about user privacy, latency, costs, and internet access requirements. In our work, we evaluate the local inference performance of various LLMs and propose a prototype capable of running LLM inference entirely on-device.
@@ 3 Research Methods and Materials
This section outlines our research methodology, including the key challenges, the LLM deployment approach, and the evaluation metrics and models used in this study.
Problem Statement: Given a set of LLMs =$\mathcal{M}=\{m_{i}\}_{i=1}^{N}$ and a set of XR devices $\mathcal{D}=\{d_{j}\}_{j=1}^{K}$, AIvaluateXR evaluates the performance of each model-device pair $(m_{i},d_{j})$ across several key factors and quantitative metrics. It first generates evaluation results for each metric, which are then passed to a Pareto analysis module to determine the Pareto-optimal set. We used AIvaluateXR to deploy 17 LLMs, denoted as $\{m_{1},m_{2},\dots,m_{17}\}$ (see Table I), across four XR and mobile devices: $d_{1},d_{2},d_{3},$ and $d_{4}$, representing Magic Leap 2 (ML2), [^3] https://www.magicleap.com/magicleap-2 Meta Quest 3 (MQ3), [^4] https://www.meta.com/quest Vivo X100 Pro (Vivo), [^5] https://www.vivo.com/en and Apple Vision Pro (AVP), [^6] https://www.apple.com/vision-pro respectively. Our goal is to perform a comprehensive evaluation of these models on the four devices to determine how well they handle large-scale language processing tasks in resource-constrained XR environments.
We define our evaluation using six key performance metrics: $p_{0},p_{1},p_{2},p_{3},p_{4},p_{5}$, which represent Model Quality, Performance Consistency, Processing Speed (w.r.t. string length and prompt length), Parallelism (w.r.t. thread count and varying batch size), Memory Usage, and Battery Usage, respectively. Fig. 1 presents an overview of our research pipeline, where each rectangular box represents a specific procedure, and the connected ovals indicate their respective outcomes. These are discussed in detail below.
@@ 3.1 LLM Deployment on XR Devices
Figure 1: Overview of our research process. The rectangular boxes represent the procedures, while the ovals indicate the outcomes. We deployed 17 LLMs on four XR devices. First, we identified the key factors affecting performance. Based on these factors, we defined our evaluation criteria and conducted the experiments. Finally, we applied 3D Pareto analysis, which provided the Pareto front to highlight the best-performing model-device pairs. We also analyzed on-device LLMs versus cloud-based LLMs (highlighted in pink).
@@ 3.3 Evaluation Approach
Considering the key challenges and our initial investigation (see Section 3.2), we formulate a standardized evaluation approach for our assessments. First, we designed an experimental setup to minimize potential biases and ensure the validity of the results (see Section 4). This setup establishes identical parameters across all devices and defines the test types to be executed. We developed several guidelines to ensure a comprehensive and standardized evaluation. To account for potential performance variations among LLMs, we selected 17 models from five different series (see Table I) and evaluated them across four devices. Additionally, we conducted Model Quality analysis to provide technical insights into the selected models. These tests are device-independent, and were conducted to understand the details and properties of each model. Next, each model-device pair was subjected to various tests (see Section 3.4).
@@ 3.4 Models and Evaluation Metrics
We use 17 LLMs from five different series (see Table I), obtained from Hugging Face’s model repository [47] in our evaluations. For consistency in naming conventions, we assigned unique IDs ($m_{1}$ to $m_{17}$) to these models.
@@ 4 Experimental Setup
Processing Speed: In this study, we measure the processing speed for each model-device pair $(m_{i},d_{j})$ by recording the time $t_{ijk}$ taken by model $m_{i}$ running on device $d_{j}$ to process a string of length $L$. The instantaneous processing speed, $S_{inst.}$, expressed in tokens per second, is calculated as: $S_{inst.}=\frac{L}{t_{ij}}$. However, instead of relying on a single run, we perform 5 runs for each model-device pair to ensure accuracy and measure stability. If the coefficient of variance (CV) across the 5 runs is less than $33\%$, we calculate the 5-Run-Mean Speed (referred to as consistent speed, $S_{c}$) as:
@@
where $S_{\text{inst.}}(k)$ represents the instantaneous speed during the $k^{\text{th}}$ run of each experiment. This approach minimizes the impact of factors such as initialization delays, hardware fluctuations, or other environmental inconsistencies, providing a more reliable measure of processing speed.
@@
We conducted the following experiments.